# New Coevolution Dynamic as an Optimization Strategy in Group Problem Solving

### v1

@@ Abstract
In society, it is common to face problems that require collaboration with other people, from everyday challenges to complex tasks, such as group projects at work. In this context, the search for more effective problem-solving strategies becomes a topic of great interest. This paper presents a new dynamic that integrates coevolution into the NK model together with an Erdös-Rényi random network, allowing more than one neighbor update. We explore how this coevolution can be employed as an optimization strategy for solving group problems. In recent coevolution models, only one agent is removed and another is added to the neighborhood of influence. Here, we allow $L$ agents to be changed, which had a large impact on the system performance. In the analysis of the results, we consider rewiring as a way for the target agent to obtain information from individuals or groups that were not part of its neighborhood. Our simulations demonstrate that coevolution can produces gain on the computational cost.
@@ 1 Introduction
Social interactions are inherently complex such that in recent years numerous models of social behavior using Network Science have emerged to explore this complexity [1]. Each new model provides a fresh perspective on human interaction [2]. Presently, a prominent aspect under exploration within complex network models is coevolution. In many complex systems the dynamic processes that occur in a network are interconnected with changes in the network topology [3, 4]. The rationale behind this focus is that co-evolutionary, or adaptive, network models offer a more accurate portrayal of real world systems compared to static or evolving networks. Empirical networks display both network dynamics (the evolution of network topology) and node state dynamics. Therefore, coevolution can be conceptualized as two simultaneous processes – link rewiring and state dynamics – mutually influencing each other [5, 6, 7]. A network is composed of a set of nodes (where each node represents an element of the system) connected by links (where each link represents the interaction between the nodes). Despite being simple, this structure is capable of encoding information present in a wide variety of natural, artificial and social interaction systems [8, 9].
@@
When considering cooperative processes, it is acknowledged that analyzing a group performance in problem-solving can be intricate [10]. Initially, one might assume that a larger number of individuals working towards the same objective would expedite reaching a solution. However, interpersonal interactions (such as everyone engaging in discussion) can lead to suboptimal outcomes, deviating from the desired solution. Different types of networks are found in the literature (square network, RGG, Erdös and Rényi, among others) and the choice of each one depends on the type of problem one wants to address. This is because the network determines the model communication pattern: who interacts with whom. The model determines what happens when someone interacts with someone.
Given the way our society is structured, human cooperation is crucial for social and technological development. We can understand this cooperation as interactions between peers and coordinated behavior in larger groups [11]. From a social perspective, each node in the network can represent an individual or a group of individuals, and each link represents the connection with people or groups that have a connection with the target agent. In this study, we consider rewiring as a way for the target agent to obtain information from individuals or groups that were not previously part of its neighborhood.
In this work we present a new dynamics with coevolution considering an Erdös-Rényi random network and we use the NK model [12]. Usually in coevolution models only one agent is removed and another one is added to the influence neighborhood. Here we allow $L$ agents to be altered, which had great impact on the performence of the system. This work is organized as follows: In Section II we present a brief discussion about the NK model and the construction of reliefs. In Section III we briefly discuss the Erdös-Rényi random network and present the proposed new co-evolution dynamics. Section IV is dedicated to presenting the results obtained. Section V presents our final considerations.
@@ 2.1 NK Fitness Landscapes
[…]
@@
For $K=0$ the function $f(\phi)$ has only one maximum $f(\phi_{M})=f_{M}$, the global one. Therefore the difference $f(\phi_{M})-f(\phi)$ is proportional to the Hamming distance $d_{H}(\phi_{M},\phi)$ between the vectors. The Hamming distance gives the number of different bits between the two vectors. For example $d_{H}(\phi_{1},\phi_{2})=1$ for $\phi_{1}=(010101)$ and $\phi_{2}=(000101)$. With $K>0$ the problem becomes more complicated because the function $f$ has additional local maxima beside the global one. Thus the difference $f(\phi_{m})-f(\phi)$ is no longer proportional to $d_{H}(\phi_{m},\phi)$. The average number of local maxima increases with $K$ until the limiting case of $K=N-1$, in which the landscape becomes completely random, known as the Random Energy model [12].
The dynamics of the temporal evolution begins at $t=0$ with the selection of an agent at random for a string update. We shall call this selected agent as the target one. This update can happen by two different ways: copy or mutation. The mutation happens when the target agent has no neighbors or when it is the model of its influence neighborhood (which is the set of neighbors agents of the target). Been the model means that the target agent has the highest fitness of his neighborhood. On the other hand, the copy will happen when the target has neighbors and it is not the model. Both processes change the string of the target agent. The mutation is achieved by a flip of a random bit of the string. The copy process involves more steps: the different bits bewteen the target and the model strings are obtained. Then, one of them is randomly selected and flipped on the target string. As a result, the target string gets more similar to the model string.
@@
The efficiency of coevolution is measured by the total number of iterations $t_{g}$ required by the first agent who finds the global maximum. A smaller $t_{g}$ means more efficient search: the goal was achieved in less time. One iteration refers to $N$ updates of $N$ randomly selected agents. The normalized computational cost is then defined as $C\equiv Mt_{g}/2^{N}$, so it can be compared between systems of different size $M$.
The independent search consider $M$ agents and the updates are only mutation. This case is used as an reference and has an analytic result [14]:
$\left\langle C\right\rangle=\frac{M}{2^{N}\left[1-\left(\lambda_{N}\right)^{M}\right]},$ (1)
where $\lambda_{N}$ depends on the value of $N$. From the definition of $C$ and Eq. 1, the expected number of trials for a group of $M$ independente agents to find the global maximum is:
$\left\langle t_{g}\right\rangle=\frac{1}{\left[1-\left(\lambda_{N}\right)^{M}\right]}.$ (2)
@@ 2.2 Random Network
A network or a graph comprises a collection of vertices interconnected by a set of edges [15]. Each edge, or connection, links a pair of vertices with the number of vertices defining the network size and the set of edges determining its connectivity. Vertices that share an edge are termed neighbors. Throughout this work the terms vertices, agents, and individuals have the same meaning.
The degree $k_{i}$ of a vertex refers to the number of connections it establishes with other vertices while the network connectivity is characterized by the distribution $P(k)$. Various types of networks are used for solutions of physical and biological problems. In this work we will use the Erdös-Rényi (ER) network [16, 17, 18, 19] which has the following definition: consider a set of $M$ vertices and a parameter $\beta$ such that $1.00.0<\beta<$ 1.0, the probability of any two vertices being connected is $\beta$. The average degree $\mu$ of the network is defined as $\mu=(1/M)\sum_{i}^{M}k_{i}$ and in the case of ER network it is $\mu=\beta(N-1)$ [20]. So we use $\mu$ as control parameter of the network instead of $\beta$.
2.3 Coevolution
The general idea of coevolution is to apply a rewiring before each string update: removing a random neighbor and adding a new one to the target agent [5]. However, we can let the number of destroyed and created connections be greater than one. Therefore, before each string update we remove $L<k_{i}$ random neighbors and add $L$ new random neighbors to the target agent. The parameter $L$ is an input to our model, it is the control parameter of the coevolution process.
The rewiring itself is the temporal evolution of the network. To be considered as coevolution, this network evolution must be coupled to the evolution of the agent string. The constant rewiring (always happening) is not coevolution, although is a network evolution. The no coevolution case is $L=0$. We have tested several different mechanisms to implement this coupling and found one which gave interesting results. The coevolution happens with a probability given by $r^{q}$ where:
$r=1-\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}},$
where $\phi_{t}$ is the fitness of the target agent, $\phi_{s}$ is the smallest fitness and $\phi_{m}$ is the maximum one (fitness of the model agent) on the influence neighborhood. With this definition, if the target is the model agent we have $r=0.0$. If it has the lowest fitness $\phi_{s}$ we have $r=1.0$. Otherwise we have $1.00.0<r<$ 1.0. The coevolution ($q>0$) is a selective rewiring, in contrast with the constant rewiring case when $q=0$. When the coevolution happens, $L$ neighbors of the target agent are removed and new $L$ random ones are created. The parameter $q$ is the non-linear one: $q=1$ is the linear case, $q=0$ means constant rewiring (always happening) and otherwise is the non-linear case. This parameter $q$ is added to the rewiring probability definition to allow studies of possible non-linear effects, as observed by for example by Raducha et al [5].
@@ 3 Results
A schematic illustration for this new dynamic is presented in Figure 1. The steps of one iteration are:
  • At time $t=0$ we initialize the network and randomly select a node to serve as the target agent.
  • The rewiring happens with probability of $r^{q}$.
  • The string update can be either a mutation or a copy. Mutation arises when the target agent is the model one. Otherwise copying takes place with probability $p$ and mutation with probability $1-p$.
$N$ iteractions are one Monte Carlo Step (MCS). The simulation concludes when the target agent achieves the highest global fitness after $t_{g}$ MCS, which is one time evolution. The computational cost $\left\langle C\right\rangle$ is averaged over $S$ independent time evolutions, which one with a random sample of the network.
Figure 1: Schematic illustration of the coevolution dynamics used. The initial network is the left bottom one and the final network is the center top one. The colors represent different components. In the initial network the target node is enclosed within a black dashed rectangle. $r_{1}$ and $r_{2}$ are random numbers from a uniform distribution between 0 and 1. The parameter $q$ is the non-linear one: $q=1$ is the linear case, $q=0$ means constant rewiring (always happening) and otherwise is the non-linear case.
In this work we explore two values for the $K$ parameter: $K=0$ is the easy case and $K=4$ is the complicated one. We study the linear case $q=1$ except when it is indicated otherwise. We also fixed $N=12$ and the copy probability at $p=0.5$. At this value of $N$ we have $\lambda_{12}\approx 0.99978$ and $t_{1}=\left\langle t_{g}\right\rangle\approx 4545$ for a system with only one agent (from Eq. 2). Since $(\lambda_{12})^{M}\approx e^{-M(1-\lambda_{12})}$, we can write the approximated results [12]:
$\left\langle C\right\rangle\approx\left\{\begin{array}[]{rl}t_{1}/2^{N},&\text{ for }M<<t_{1}\\ M/2^{N},&\text{ for }M>>t_{1}\end{array}\right.$
At $N=12$, the limit value for low $M$ is $\left\langle C\right\rangle=t_{1}/2^{12}\approx 1.110$. We compare our results with 3 others data sets:
  • Independent Search (IS): when each agent independently searches for the global maximum. In this case there is no copy of the model strings, only mutation. It can be achieve by setting $p=0$ or using a network with $k_{i}=0$ for all agents $i$. This result will be displayed as a green solid line on the results.

### v2

@@ Abstract
Coevolution on social models couples the time evolution of the network with the time evolution of the states of the agents. This paper presents a new coevolution dynamic allowing more than one rewiring on the network. We explore how this coevolution can be employed as an optimization strategy for problem-solving capability of task-forces. We used an agent-based model study how this new coevolution dynamic can help a group of agents whose task is to find the global maxima of NK fitness landscapes. Each agent can replace more than one neighbors, and this quantity is a tunable parameter in the model. These rewirings is a way for the agent to obtain information from individuals that were not previously part of its neighborhood. Our results showed that this tunable coevolution can indeed produce gain on the computational cost under certain circunstances. At low average degree (using a random network to connect the agents) and on easy landscape, more rewirings return lower cost, helping the agents to find the global maximum faster. However, at larger average degree and difficult landscape, the effect is more complex. At medium size systems, the coevolution brings optimal results when 3 or 4 neighbors are replaced.
@@ 1 Introduction
In society, it is common to face problems that require collaboration with other people, from everyday challenges to complex tasks, such as group projects at work. In this context, the search for more effective problem-solving strategies becomes a topic of great interest.
Social interactions are inherently complex such that, in recent years, numerous models of social behavior have emerged to explore this intricacy [1, 2, 3] using the framework of Network Science [4]. The nodes of the network represent individuals (or agents) with states defined by the social model while the connections represent the interactions between different individuals. The states of the individuals evolve following the set of rules from the dynamic social model in study. The network itself can also have a time evolution: links can be destroyed or created.
@@
Presently, a prominent aspect under exploration within dynamic social processes on complex networks is coevolution: the evolution of the agent states is coupled to the evolution of the network [5, 6, 7]. This coevolution has been studied in different social models [8, 9, 10, 11, 12, 13, 14] and Evolutionary Games Theory [15, 16]. The rationale behind this dynamics is that co-evolutionary, or adaptive, network models offer a more accurate portrayal of real world systems compared to static or evolving networks. Empirical networks display both network dynamics (the evolution of network topology) and node state dynamics. Therefore, coevolution can be conceptualized as two simultaneous processes – link rewiring and state dynamics – mutually influencing each other
The central idea of coevolution in social dynamics is the agent uses the social information (states of the agents) of his influence neighborhood (set of neighbors) to remove one neighbor and add a new one. And then, considering this new influence neighborhood, his social state will evolve in time accordingly. This mechanism can make a difference in cooperative processes, where one studies the performance of a set of individuals solving a given task [17, 18, 19]. Each individual not only use his influence network, but also changes it if he thinks one neighbor may not give a proper contribution to solve the problem. This is specially important in imitative learning, where one tries to learn from an expert’s experience and behavior [20, 21].
@@
How to modify the influence network is an important point to achieve success. Analyzing a group performance in problem-solving can be intricate. Initially, one might assume that a larger number of individuals working towards the same objective would expedite reaching a solution. However, interpersonal interactions (such as everyone engaging in discussion) can lead to suboptimal outcomes, deviating from the desired solution [22, 23].
In this work we study the effect of a coevolution mechanism on a cooperative process problem. To address this question, we use an agent-based model where the agents can perform individual trial-and-test searches to probe a fitness landscape (exploration) or imitate a model agent – the best performing agent in their influence neighborhood at the trial (exploitation). The coevolution mechanism gives a chance for the agent to change his influence neighborhood on the exploitation case. We consider a scenario where the agents are fixed at the nodes of a random network and can interact with each other if they are connected. In addition, before each interaction the agent has a chance to change $L$ agents within his influence neighborhood. This set of rules creates a time-dependent and adaptive network.
This work is organized as follows: In Section 2.1 we present a brief discussion about the NK model, in section 2.2 we define the computational cost, in section 2.3 we define how the agents are updated, in section 2.4 we define the proposed coevolution mechanism and in section 2.6 we described the random network used to connect the agents. Section 3 is dedicated to present and discuss the results and section 4 is dedicated to our final considerations.
@@ 2.1 NK Fitness Landscapes
The NK model was introduced by Kauffman and Levin as a fitness model for adaptive evolution processes such as walking in rugged landscapes [24, 25]. This model has been extensively used in different studies of cooperative process and imitative learning [26, 27, 28, 29, 30]. It is well suited for example for the imperfect version of the imititive learning process [31], when one agent copies only part of the solution from an expert.
The system is composed of $M$ agents described by a binary string (or vector) of $N$ bits $\mathcal{X}=(x_{1},x_{2},...,x_{N})$ with $x_{i}\in\{0,1\}$. In total there are $2^{N}$ different vectors $\mathcal{X}$. Each string has a fitness $f(\mathcal{X})$, limited in the interval $1.00.0<f(\mathcal{X})<$ 1.0, calculated as [32]:
$f(\mathcal{X})=\frac{1}{N}\sum_{i=1}^{N}\Phi_{i}(K,\mathcal{X}),$ (1)
where $\Phi_{i}(\mathcal{X})$ is the contribution of component $i$ to the fitness of string $\mathcal{X}$. It depends on the state $x_{i}$ as well as on the states of the $K$ right neighbors of $i$, i.e.,
$\Phi_{i}(K,\mathcal{X})=\Phi_{i}(x_{i},x_{i+1},x_{i+2},...,x_{i+K}),$ (2)
with the arithmetic in the subscripts done modulo $N$ [33]. The functions $\Phi_{i}$ are $N$ distinct real-valued functions on $\left\{0,1\right\}^{K+1}$. As usually done, we assign to each $\Phi_{i}$ a uniformly distributed random number in the unit interval, so that $f(\mathcal{X})\in(0,1)$ [34].
@@
The function $f(\mathcal{X})$ has a global maximum $f(\mathcal{X}_{M})=F$ at the string $\mathcal{X}_{M}$. The number of local maxima of the function $f(\mathcal{X})$ is zero at $K=0$ landscape. At $K>0$ the number of local maxima increases. Therefore, at $K=0$, the difference $f(\mathcal{X}_{M})-f(\mathcal{X})$ is proportional to the Hamming distance $d_{H}(\mathcal{X}_{M},\mathcal{X})$ between the vectors [35]. The Hamming distance gives the number of different bits between the two vectors. For example, consider $\mathcal{X}_{1}=(010101)$ and $\mathcal{X}_{2}=(000101)$. The hamming distance is $d_{H}(\mathcal{X}_{1},\mathcal{X}_{2})=1$. At $K>0$, the landscape becomes more complicated because the function $f$ has additional local maxima beside the global one. Thus the difference $f(\mathcal{X}_{M})-f(\mathcal{X})$ is no longer proportional to the hamming distance $d_{H}$. The average number of local maxima increases with $K$ until the limiting case of $K=N-1$, in which the landscape becomes completely random, known as the Random Energy model [36].
2.2 Computational cost
The time evolution of the system is measured in Monte Carlo steps (MCS). The evolution gets to a halt when one agent reaches the maximum fitness $F$ of the landscape. Each agent changes its fitness $f(\mathcal{X})$ when its string $\mathcal{X}$ is updated (described in next section). So, all agents try to increase their fitness during the time evolution. The dynamics begins at $t=0$ with the selection of one agent for a string update. One Monte Carlo step (MCS) is the update of $M$ random agents, in sequence.
@@ 2.2 Computational cost
The efficiency of the search is measured by the time $t_{g}$ in MCS required by the first agent to find the global maximum $F$, in other words, when its string becomes $\mathcal{X}_{M}$ after a number of updates. A small $t_{g}$ means more efficient search: the goal was achieved in less time. So, the lower the better. The normalized computational cost is then defined as [11]:
$C\equiv Mt_{g}/2^{N}.$ (3)
This normalization ensures $C$ can be compared between systems of different sizes $M$ and landscapes $N$.
It is important to compare this computational cost with the one from the system considering $M$ independet agents. This means that each agent does not interact with any other. In this case, the agents are placed on the nodes of a completely disconnected network, containing $M$ components. This case is used as an reference and has an analytic result [37]:
$\left\langle C_{I}\right\rangle=\frac{M}{2^{N}\left[1-\left(\lambda_{N}\right)^{M}\right]},$ (4)
where $\lambda_{N}$ depends on the value of $N$. $C>C_{I}$ means the interaction between the agents creates more problems than solution, slowing down the quest of the agents for the global maximum. Otherwise, if $C<C_{I}$, the interaction helped the agents making one of them find the global maximum in less time.
2.3 Strings update
Each agent has its string $\mathcal{X}$ updated when it is selected for analysis. This is how each agent tries to improve its fitness $f(\mathcal{X})$, to eventually reaches the global maximum. We call this agent the target, and he can implement one of the two strategies to perform the update: exploitation or exploration [33]. Exploitation implies the target tries to learn the solution from a neighbour with a higher fitness. This is implemented as a copy, with probability $p$: the target copies one bit of the string of the model agent [35]. Otherwise, exploration means the target aims to innovate a new solution with a better fitness by itself. It is a mutation on the string of the target, and it should happen with probability $1-p$. However if the target has no neighbors or if it is the model of its influence neighborhood, the mutation will happen for sure. If it has no neighbors, it does not have anyone to copy. And if it is already the model, there is no point in copy another agent with a smaller (worst) fitness. The model agent is the one with the highest fitness in the target influence neighborhood. The target agent may be the model if it has a larger fitness than his neighbors.
Both processes change the string of the target agent. The mutation is achieved by a flip of a random bit of the string. The copy process involves more steps: the different bits bewteen the target and the model strings are selected. Then, one of them is randomly selected and flipped on the target string. This is the imperfect copy on the imitation learning context [31]. As a result, the target string gets more similar to the model string (hamming distance decreases). If $K=0$, this means the target agent increases his fitness. On the other hand, if $K>0$, the model one may be a local maxima, and the hamming distance between the target string $\mathcal{X}$ and the global maximum one $\mathcal{X}_{M}$ may increase when the target copies the local maximum (model agent). The target can get trapped on the suboptimal solution (local maximum) [18]. $K=0$ is the easy landscape and increasing $K$ increases the difficulty of the landscape. The most difficult landscape is $K=N-1$ where there is no point on copy because the landscape is completely random.
2.4 Coevolution mechanism
The general idea of coevolution is to apply a rewiring before each string update: removing a random neighbor and adding a new one to the target agent. However, we can let the number of destroyed and created connections be greater than one. Therefore, before each string update we remove $L<k_{i}$ random neighbors and add $L$ new random neighbors to the target agent. The parameter $L$ is an input to our model, it is the control parameter of the coevolution process. The static network scenario is recovered with $L=0$.
Figure 1: Schematic illustration of the implemented model. (a) Illustration of one iteration. It starts with the selection of the target agent. $r$ is calculated through Eq. 6. With probability $r^{q}$ the network can suffer a rewiring. The next step is the string update, where a copy happens with probability $p$ and a mutation with probability $1-p$. (b) Illustration of the rewiring process. The target agent (indicated by a dashed black square) loses $L$ neighbors and gets new $L$ ones. The case $L=1$ is illustrated in this picture: the target loses the neighbor on the right and gets acquainted with the agent right above him. The colors of the nodes indicate the components.
The rewiring itself is the temporal evolution of the network. To be considered as coevolution, this network evolution must be coupled to the evolution of the agent string. The constant rewiring (always happening) is not coevolution, although it is a network evolution [35]. As the rewiring itself is a random process, the coupling of the network evolution with the agent states, in our model, will be achieved by the probability of rewiring $r$: it must be calculated considering the fitness of the agents. The rewiring process happens before the string update, as illustrated on Fig. 1(a). Another illustration of one example of the rewiring itself (with $L=1$) is shown Fig. 1(b).
2.5 Rewiring probability
The goal of each agent is to increase its fitness until one of them, eventually, reaches the global maximum of the landscape. So, having neighbors with better fitness if beneficial. The probability $r$ should be larger when most of the neighbors have smaller fitness, so one of them may be replaced with a higher fitness one. The limit cases of this hypothesis are:
  • the agent is the model: all neighbors have smaller fitness, so it is justifiable to replace some of them. The probability should be maximum: $r=1$.
  • the agent has the smallest fitness of the neighborhood: any copy is already rewarding, so there is no urgency to replace any any neighbor. The probability should be minimum: $r=0$.
  • Otherwise, the probability shoud be such as $10<r<$ 1.
This behavior is accomplished by the definition:
$r=\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}},$ (5)
where $\phi_{t}$ is the fitness of the target agent, $\phi_{s}$ is the smallest fitness and $\phi_{m}$ is the maximum one (fitness of the model) on the influence neighborhood. Notwithstanding, after extensive tests, we did not found any gain on the computational cost with coevolution using this probability.
We have tested several other definitions for the probability, surprinsingly, only one gave interesting results:
$r=1-\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}}.$ (6)
This is the opposite of the first attempt: $r=0$ if the target is the model and $1$ if it has the smallest fitness. The assumption from Eq. 5 is based on the update via copy. It is worth mention that this copy happens with probability $p$, been the other update process the mutation. So, regardless Eq. 5 seems more plausible, it is possible that other mechanisms (such as Eq. 6) may yield better results. After all, the outcome after a mutation is independent if there was a rewiring or not.
For completeness sake, we add one more parameter to the probability, giving it the form $r^{q}$. The free parameter $q\geq 0$ controls the frequency of the rewiring: $q=1$ (probability is just $r$) enables the coevolution and $q=0$ (probability equals $1$) implies a constant rewiring. Possible non-linear effects can be examined using other values subjected to the constraint $q\neq 0$ and $q\neq 1$ [38].
@@ 2.6 Random Network
A network or a graph comprises a collection of vertices interconnected by a set of edges [39, 40]. Each edge, or connection, links a pair of vertices with the number of vertices defining the network size and the set of edges determining its connectivity. Vertices that share an edge are termed neighbors. Throughout this work the terms vertices, agents, and individuals have the same meaning. Various types of networks are used for solutions of physical and biological problems [41, 42, 43]. In this work we will use the Erdös-Rényi (ER) network [44, 45] which has the following definition: consider a set of $M$ vertices and a parameter $\beta$ such that $10<\beta<$ 1, the probability of any two vertices being connected is $\beta$. This network has been used frequently as a first prototype to study the influence of the network topology on social dynamics [46, 47, 48].
The degree $k_{i}$ is the number of neighbors of node $i$ and the average degree $\mu$ of the network is defined as $\mu=(1/M)\sum_{i}^{M}k_{i}$ [49]. In the case of ER network, we have $\mu=\beta(M-1)$ [4]. So we use $\mu$ as control parameter of the network instead of $\beta$.
@@ 3 Results
A schematic illustration for our dynamic is presented in Figure 1. The input parameters are $M$, $L$, $p$, $N$, $K$ and $S$. The general algorithm is:
  1. 1.
    A sample of the random network is created and the time starts at $t=0$.
  2. 2.
    $t_{g}$ Monte Carlo Steps are executed until one agent finds the landscape maximum $F$. The computational cost is calculated (Eq. 3).
  3. 3.
    Steps 1 and 2 constitute one time evolution. $S$ random time evolutions are perfomed and the average $\left\langle C\right\rangle$ computational cost is recorded.
One Monte Carlo step is the execution of $M$ iterations. The algorithm for one iteration is:
  • One random node is selected to serve as the target agent.
  • The rewiring happens with probability of $r^{q}$ where $L$ links are removed and new $L$ links are created involving the target agent.
  • The string update takes place and it can be either a mutation or a copy.
Fig. 1(a) illustrates this algorithm and 1(b) shows an example of the rewiring with $L=1$.
Figure 2: Results of the Reference Case on the two investigated landscapes. Computational cost $\left\langle C\right\rangle$ as function of $M$ for different average degrees $\mu$ of the ERN with no coevolution. These results is the reference to which we will compare the results with coevolution to evaluate its impact. Parameter: $q=1$ and $S=5\times 10^{4}$ samples. The green line is the IS (Eq. 4), black dashed line is the FCN result and the pink line is the approximation for large $M$. Purple, blue, green, orange and red circles correspond to an average degree of $\mu=1$, 5, 10, 20 and 100, respectively. (a) Easy landscape: $K=0$. (b) Difficult landscape: $K=4$.
Figure 3: Snaphots of instances of the random network used in this work. System size: $M=67$ agents. Each component is represented in a different color. (a) Low average degree scenario: $\mu=1.0$. The network is not percolated and the largest component is in painted in red. (b) High average degree scenario: $\mu=10.0$. In this scenario the network is percolated.
In this work we explored two values for the $K$ parameter: $K=0$ is the easy case and $K=4$ is the complicated one. We also investigated the following values for the coevolution: $L=0$, 1, 2, 3, 4, 5 and 6. We also fixed $N=12$ and the copy probability at $p=0.5$. At this value of $N$ we have $\lambda_{12}\approx 0.99978$. Since $(\lambda_{12})^{M}\approx e^{-M(1-\lambda_{12})}$, we can approximate the limit cases (low and large $M$) of the independent search from Eq. 4 as [28, 37]:
$\left\langle C_{I}\right\rangle\approx\left\{\begin{array}[]{rl}1.110,&\text{ for }M<<4545\\ M/2^{12},&\text{ for }M>>4545\end{array}\right.$ (7)
3.1 Benchmark
In order to better assess the impact of the coevolution porcess, we compare its results to 3 others data sets:
  • Independent Search (IS): when each agent independently searches for the global maximum. In this case there is no copy of the model strings, only mutation. Numerically, it can be achieved by setting $p=0$ or using a network with $k_{i}=0$ for all agents $i$. Analytically, it can be calculated using Eq. 4, and its limit cases for small and large $M$ are given by Eq. 7. This result will be displayed as a green solid line on the results.